OpenStudy (anonymous):

-5 x Y + 15= -125 I know the answer but I don't know how to get the steps to the answer.

OpenStudy (anonymous):

Where you have -5 x Y, is the "X" supposed to be for multipulcation?

OpenStudy (anonymous):

Yes it's multiplication Niall

OpenStudy (anonymous):

-5 x Y + 15 = -125 -5Y = -125 - 15 -5Y = -140 Y = -140/-5 Y = 28

OpenStudy (anonymous):

Thank you so much sheevani.

OpenStudy (anonymous):

Of course! :)